Check my work Required information [The following information applies to the Bargain Rental Car offers rental cars in an off-airport location near a major tourist destination in Californ would like to better understand the variable and fixed portions of it car washing costs. The company operates its wash facility in which each rental car that is returned is thorough customer. Management bel Accordingly, the following data have been compiled: own car before being released for rental to another lieves that the variable portion of its car washing costs relates to the number of rental returns r Wash Costs January February March April Hay June uly August September October November December 2,480 2,589 2,880 3,800 3,780 5,100 5,5e0 5,50e 4,780 4,180 2,200 2,908 11,206e 13,200 12,00e s 14,600 16,400 s 23,700 s 22,480 s 22,780 23,000 s 21,700 5 18,908 s 14,900 2. Using least-squares regression, estimate the variable cost per rental return and the monthly fixed cost incurred to wash cars. (Round Fixed cost to the nearest whole dollar amount and the Variable cost per unit to 2 decimal places.)
Answer #1

2. Variable cost per rental return = $3.92

Month Rental Returns Car wash costs
x y x^2 x*y
January 2400 11200 5760000 26880000
February 2500 13200 6250000 33000000
March 2800 12000 7840000 33600000
April 3000 14600 9000000 43800000
May 3700 16400 13690000 60680000
June 5100 23700 26010000 120870000
July 5500 22400 30250000 123200000
August 5500 22700 30250000 124850000
September 4700 23000 22090000 108100000
October 4100 21700 16810000 88970000
November 2200 10900 4840000 23980000
December 2900 14900 8410000 43210000
Total 44400 206700 181200000 831140000
?x = 44,400
?y = 206,700
?xy = 831,140,000
?x^2 = 181,200,000
Unit variable cost = b = (n*?xy - ?x*?y)/(n?x^2 - (?x)^2)
=(12*831,140,000 - 44,400*206,700) /
                      (12*181,200,000 - 44,400*44,000)
=796,200,000 / 203,040,000
                             =3.92
Total Fixed Cost = a = (?y -b*?x)/n
=(206,700 - 3.92*44,400)/12
                             =2,852